# Social Bias in Large Language Models For Bangla: An Empirical Study on Gender and Religious Bias

### v1

@@ Abstract
The rapid growth of Large Language Models (LLMs) has put forward the study of biases as a crucial field. It is important to assess the influence of different types of biases embedded in LLMs to ensure fair use in sensitive fields. Although there have been extensive works on bias assessment in English, such efforts are rare and scarce for a major language like Bangla. In this work, we examine two types of social biases in LLM generated outputs for Bangla language. Our main contributions in this work are: (1) bias studies on two different social biases for Bangla (2) a curated dataset for bias measurement benchmarking (3) two different probing techniques for bias detection in the context of Bangla. This is the first work of such kind involving bias assessment of LLMs for Bangla to the best of our knowledge. All our code and resources are publicly available for the progress of bias related research in Bangla NLP.  [^1] https://github.com/csebuetnlp/BanglaSocialBias
@@ 1 Introduction
The rapid advancement of Large Language Models (LLMs) has significantly impacted various domains, particularly in social influence and the technology industry Kasneci et al. (2023); Dong et al. (2024b). Given their growing influence, it is crucial to ensure LLMs are free from harmful biases to avoid legal and ethical issues Weidinger et al. (2021); Deshpande et al. (2023). In the context of computing systems, bias is where sociotechnical systems systematically and unfairly discriminate against certain individuals or social groups in favor of others Friedman and Nissenbaum (1996); Blodgett et al. (2020). Hence, analyzing bias and stereotypical behavior in LLMs is vital for identifying and mitigating existing biases, thereby fostering the development of responsible and ethical models.
Bangla, the sixth most spoken language globally with over 230 million native speakers constituting 3% of the world’s population [^2] https://w.wiki/Psq, is underrepresented in NLP literature due to a lack of quality datasets (Joshi et al., 2020). This gap limits our understanding of bias characteristics in language models, including LLMs. The need to measure bias in Bangla arises from this gap. Historically, societal views in Bangla-speaking regions have undervalued women, leading to employment and opportunity discrimination (Jain et al., 2021; Tarannum, 2019). Additionally, the region’s significant religious diversity, primarily among Muslims and Hindus, makes Bangla a valuable case study for examining gender and religious biases, two important social biases.
In this study, we pose the question, does multilingual LLMs exhibit gender and religious bias when prompted with Bangla?. To address this, we present: (1) a curated dataset specifically designed to detect gender and religious biases in Bangla, (2) thorough bias probing analysis on both popular and state-of-the-art closed and open-source LLMs, and (3) an empirical study on bias through LLM-generated responses.
Our findings reveal significant biases in LLMs for the Bangla language and highlight shortcomings in their generative power, underscoring the need for future de-biasing efforts.
@@ 2 Related Work
Several works on bias measurements in language models are done in recent years Mehrabi et al. (2021). Existence of gender bias has been exposed in tasks like Natural Language Understanding Bolukbasi et al. (2016); Gupta et al. (2022); Stanczak and Augenstein (2021) and Natural Language Generation Sheng et al. (2019); Lucy and Bamman (2021); Huang et al. (2021) .Benchmarks such as WinoBiasZhao et al. (2018) and WinogenderRudinger et al. (2018) have been used to measure gender biases in LMs. Preliminary studies on religious and ethnic biases were done in some works. (Milios and BehnamGhader, 2022; Navigli et al., 2023; Abid et al., 2021). Works like Nadeem et al. (2020); Nangia et al. (2020) provide frameworks and datasets for different types of biases including gender and religion.
Recent studies have conducted experiments on determining gender stereotypes in LLMs Kotek et al. (2023); Ranaldi et al. (2023); Dong et al. (2024a) and debiasing techniques Gallegos et al. (2024); Ranaldi et al. (2023), but most of them are on English. There are a few works on multilingual settings Zhao et al. (2024); Vashishtha et al. (2023), but such efforts are not common for Bangla. The most preliminary work on Bangla bias detection is found in the works of Sadhu et al. (2024), that includes static and contextualized embeddings. Notable works on bias evaluation dataset creation are done in Das et al. (2023).
@@ 3 Data
We use two strategies for LLM probing: Template-based and Naturally sourced. The template-based approach uses curated templates for gendered persona or religious group predictions to reveal biases in LLMs. Naturally sourced sentences, on the other hand, are used to make explicit predictions about groups or genders, helping to understand the LLM’s ability to interpret natural scenarios.
Template Based: We create semantically bleached templates with placeholders for specific traits that are filled with adjective words to create prompts. We use these prompts for gendered role prediction or religious identity prediction. The adjective words encompass different categories used to describe human nature e.g. Personality Based, Outlook Based and Communal Based. We also include Occupation as a category to understand the association between gender and religion with occupation. We create templates to intermix human professions with traits to study if the addition of occupation has any effect on role prediction. We were careful to not induce any stereotypes into the templates and make sure all the adjective words and occupations are equally likely for any gender or religious community. In total, we have 2772 template sentences by combining both the categories (see appendix A.1 for more details).
Naturally Sourced: We use the BIBED dataset Das et al. (2023), specifically the data from Explicit Bias Evaluation(EBE) as a source of naturally occurring scenarios. The sentences are structured in pairs, each containing one identifying subject from a group of either male-female words(for gender) or Hindu-Muslim words (for religion). We replace the main identity dimension in the sentence with _ (gap) and give the model options to select between the possible identities. We provide the examples in appendix A.2. However, there are many instances where one a sentence without the subject is not equally probable for both the contrasting identities. In order to curate sentences that serve our cause, we manually select these sentences to provide equal opportunity. Details of the selection process is given in appendix A.3. After the curation process, we are left with 2600 pairs for gender and 1627 pairs for religion.
Table 1: The prompt template and an example of prompt for gender role prediction (Note that the translations are only for understanding and not used in prompting)
@@ 4.1 Model Selection
For our experiment we provide results for three state-of-the-art LLMs: Llama3-8b (version: Meta-Llama-3-8B-Instruct 333meta-llama/Meta-Llama-3-8B-Instruct) (AI@Meta, 2024), GPT-3.5-Turbo 444gpt-3-5-turbo and GPT-4o 555gpt-4o. Since Bangla is a low resource language, not many models could generate the expected response we required. For our experimentation, we also tried some other models (mentioned in limitations) but none could produce any presentable result that serves our purpose.
@@ 4.2 Prompt
We design prompts to best fit the generation task. In the case of template based probing, we prompt the model for gendered role or religious identity selection, and in the case of naturally sourced probing, we use fill in the blanks approach. We provide an example of prompt creation for gender in template based probing in table 1.
Since the task of token prediction and generation is a stochastic process, we repeat each prompt twice to bring stable result. In table 2 we provide the number of prompts for each model.
@@
Template Based Gender 4256
Religion 1288
Naturally Sourced Gender 5200
Religion 3254
@@ 4.3 Evaluation Setup
Evaluation Metric: We use a widely used fairness measurement metric, Disparate Impact(DI) Feldman et al. (2015) for evaluation. It is computed by $\frac{P(Y=1|S\neq 1)}{P(Y=1|S=1)}$ The identifiers of our experiments are all binary (e.g: male-female, hindu-muslim), so we can apply this to our cause with empirical estimation. In task Q, for a category a, where the possible outcomes are x, y, the DI calculation is
@@
The empirical estimation we are using is
$DI_{Q}(a)=\frac{C_{x}(a)}{C_{y}(a)}$
where $C_{z}$ represents the occurrence frequency of z class elements. For our experiments, we consider $DI_{G}$ and $DI_{R}$ for gender and religion predictions and $(x=female,y=male)$ and $(x=Hindu,y=Muslim)$ for the two bias evaluations respectively. For any fair LLM, it is obvious that the DI score for any category would be close to 1.
@@ 5 Results and Evaluation
(a) DI Scores for Gender Bias(Positive Traits)
(b) DI Scores for Gender Bias(Negative Traits)
(c) DI Scores for Religious Bias(Positive Traits)
(d) DI Scores for Religious Bias(Positive Traits)
Figure 1: Bias in role selection for multiple LLMs in the case of template based probing for gender and religion data. We present positive and negative traits result separately. The upper bound is set to 3 and 4 for gender and religion respectively. The neutral line $(DI=1)$ is highlighted in all the figures.
@@ 5.1 Template Based Probing Results
We present the template based results in figure 1. We report the results based on 6 different categories and include the results for positive and negative traits separately for more nuanced variations.
Gender Bias: Our findings indicate GPT-3.5 exhibits the most gender bias, with a disparity impact score above 1 in all categories, suggesting a bias towards the feminine gender. Llama-3 shows significant bias in the opposite direction, with scores well below neutral, indicating male gender bias. GPT-4o displays less gender bias, with scores close to 1 in some instances. Adding occupation to the probing strategy generally reduces bias across most categories. GPT-3.5 moves closer to neutral for positive traits, and mostly for negative traits except Outlook+Occupation. We also see the opposite shift in some cases, like GPT-4o shifting for Personality+Occupation and Communal+Occupation in positive traits. It can also be seen that GPT-3.5 shifts highly in the opposite direction when we move from positive traits to negative traits, indicating its tendency to associate the negative traits with male gender.
Religious Bias: An interesting observation in this case is the shift of model scores from >1 to <1 when we change the category association from positive traits to negative traits. This indicates that all the models tend to associate negative traits with Muslim community and positive traits with Hindu community, a clear indication of harmful social stereotyping. Llama-3 shows high level of bias (DI > 2) in Ideology category for both traits. The addition of occupation affects the DI scores mostly for Ideology category for both traits and not so much for Outlook. Thus we don’t see much contribution of Occupation on Outlook to change any behaviour of the models.
@@
5.2 Naturally Sourced Probing Results
Gender Bias: Figure 2 reveals that GPT-4o exhibits the highest DI score among the three models, indicating a significant disparity (favoring one gender over another) in its performance. In contrast, GPT-3.5 has a DI score slightly above neutral line, showing a relatively balanced performance with minor disparities. LLaMA-3, with a DI score below neutral line, indicates a disparity that favors the opposite gender compared to GPT-4o, yet is closer to the fairness threshold than GPT-4.
Religious Bias: the DI scores for religious bias in Figure 2 are comparatively closer among all models. GPT-4o and LLaMA-3 both exhibit DI scores below the neutral line, suggesting some level of bias, though less pronounced than the gender disparity observed in GPT-4. GPT-3.5, with a DI score just above the neutral line, indicates a slightly more balanced performance in religious contexts.
This implies that while improvements are needed across all models for religious contexts, the disparities are less severe than those related to gender.
Key Take-away: We found significant bias for both gender and religion for all three models utilizing both our probing techniques.
@@ 6 Conclusion
To summarize, by conducting experiments using two different probing techniques and dataset, we investigate gender and religion bias in multilingual LLMs in the context of Bangla. Our work demonstrates the existence of bias in both categories on different degrees. This emphasizes the need for de-biasing techniques to be applied for the use of LLMs in sensitive tasks in realm of Bangla Language and developing proper linguistics nuanced and culturally aware framework for bias measurement. In future, we plan to investigate the effects of bias in downstream applications of Bangla language models and expand to other social and cultural bias areas.
@@
Our study utilized closed-source models like GPT-3.5-Turbo and GPT-4o, which present reproducibility challenges as they can be updated at any time, potentially altering responses regardless of temperature or top-p settings. We also attempted to conduct experiments with other state-of-the-art models such as Mistral-7b-Instruct 666mistralai/Mistral-7B-Instruct-v0.2 (Jiang et al., 2023), Llama-2-7b-chat-hf 777meta-llama/Llama-2-7b-chat-hf (Touvron et al., 2023), and OdiaGenAI-BanglaLlama  [^8] OdiaGenAI/odiagenAI-bengali-base-model-v1 (Parida et al., 2023). However, these efforts were hindered by frequent hallucinations and an inability to produce coherent and presentable results. This issue underscores a broader challenge: the current limitations of LLMs in processing Bangla, a low-resource language, indicating a need for more focused development and training on Bangla-specific datasets.
@@
We also acknowledge that our results may vary with different prompt templates and datasets, constraining the generalizability of our findings. Stereotypes are likely to differ based on the context of the input and instructions. Finally our techniques all utilizes binary identities(male-female, Hindu-Muslim) for the constraints on dataset and techniques used. Despite these limitations, we believe our study provides essential groundwork for further exploration of social stereotypes in the context of Bangla for LLMs.
@@
Appendix A Data Preparation
A.1 Preparing Templates

### v2

@@ Abstract
The rapid growth of Large Language Models (LLMs) has put forward the study of biases as a crucial field. It is important to assess the influence of different types of biases embedded in LLMs to ensure fair use in sensitive fields. Although there have been extensive works on bias assessment in English, such efforts are rare and scarce for a major language like Bangla. In this work, we examine two types of social biases in LLM generated outputs for Bangla language. Our main contributions in this work are: (1) bias studies on two different social biases for Bangla (2) a curated dataset for bias measurement benchmarking (3) testing two different probing techniques for bias detection in the context of Bangla. This is the first work of such kind involving bias assessment of LLMs for Bangla to the best of our knowledge. All our code and resources are publicly available for the progress of bias related research in Bangla NLP.  [^1] https://github.com/csebuetnlp/BanglaSocialBias
@@ 1 Introduction
The rapid advancement of Large Language Models (LLMs) has significantly impacted various domains, particularly in social influence and the technology industry Kasneci et al. (2023); Dong et al. (2024b). Given their growing influence, it is crucial to ensure LLMs are free from harmful biases to avoid legal and ethical issues Weidinger et al. (2022); Deshpande et al. (2023). In the context of computing/socio-technical systems, bias refers to the unfair and systematic favoritism shown towards certain individuals or social groups, often at the expense of others, resulting in discriminatory outcomes Friedman and Nissenbaum (1996); Blodgett et al. (2020). Hence, analyzing bias and stereotypical behavior in LLMs is vital for identifying and mitigating existing biases.
Bangla, the sixth most spoken language globally with over 230 million native speakers constituting 3% of the world’s population [^2] https://w.wiki/Psq, has remained underrepresented in NLP literature due to a lack of quality datasets (Joshi et al., 2020). This gap limits our understanding of bias characteristics in language models, including LLMs. Historically, societal views in Bangla-speaking regions have undervalued women, leading to employment and opportunity discrimination (Jain et al., 2021; Tarannum, 2019). Additionally, the region’s cultural and historical context between two major religions, Hindu and Muslim, makes Bangla a valuable case study for examining religious biases as well.
In this study, we pose the question, to what extent do multilingual LLMs exhibit Gender and Religious Bias in Bangla context?. To address this, we present: (1) a curated dataset specifically designed to detect gender and religious biases in Bangla, (2) detailed bias probing analysis on both popular and state-of-the-art closed and open-source LLMs, and (3) an empirical study on bias through LLM-generated responses.
Our findings reveal significant biases in LLMs for the Bangla language and highlight shortcomings in their generative power and understanding of the language, underscoring the need for future de-biasing efforts and better Bangla specific finetuning of LLMs.
Figure 1: Workflow for the creation of corpus for the experiments detailed in this study.
@@ 2 Related Work
Existence of gender bias has been exposed in tasks like Natural Language Understanding Bolukbasi et al. (2016); Gupta et al. (2022); Stanczak and Augenstein (2021) and Natural Language Generation Sheng et al. (2019); Lucy and Bamman (2021); Huang et al. (2021). Benchmarks such as WinoBias Zhao et al. (2018) and Winogender Rudinger et al. (2018) have been used to measure gender biases in LMs. Preliminary studies on religious and ethnic biases are done in some works (Milios and BehnamGhader, 2022; Navigli et al., 2023; Abid et al., 2021). Works like Nadeem et al. (2021); Nangia et al. (2020) provide frameworks and datasets for different types of biases including gender and religion. IndiBias Sahoo et al. (2024), a benchmark in Indian context, has been introduced to measure socio-cultural biases in LLMs.
Recent studies have conducted experiments on determining gender stereotypes in LLMs Kotek et al. (2023); Ranaldi et al. (2024); Jha et al. (2023); Dong et al. (2024a) and debiasing techniques Gallegos et al. (2024); Ranaldi et al. (2024), but most of them are on English. There are a few works on multilingual settings Zhao et al. (2024); Vashishtha et al. (2023), but such efforts are not common for Bangla. The most preliminary work on Bangla bias detection is found in the works of Sadhu et al. (2024), that includes static and contextualized embeddings. Effectiveness of varied probing techniques for extracting cultural variations in PLMs has been discussed in Arora et al. (2023).
@@ 3 Data
We use two strategies for LLM probing: Template Based and Naturally Sourced. The workflow of our proposed approach is illustrated in Figure 1. We explain the two techniques as follows:
Template Based: We create semantically bleached templates with placeholders for specific traits that are filled with adjective words to create prompts. We use these prompts for gendered role or religious identity prediction. The adjective words encompass different categories used to describe human nature e.g. Personality, Outlook and Communal based. We also include Occupation as a category to understand the association between gender and religion with occupation. We create templates to intermix human professions with traits to study if the addition of occupation affects role prediction. We ensured that no stereotypes were induced into the templates and that all adjective words and occupations are equally probable for any gender or religious community. In total, we have 2772 template sentences by combining both the categories (see appendix C.1 for more details).
Naturally Sourced: We use the BIBED dataset Das et al. (2023), specifically the Explicit Bias Evaluation (EBE) data for naturally occurring scenarios. The sentences are structured in pairs, each containing one identifying subject from a group of either male-female words (for gender) or Hindu-Muslim words (for religion). We replace the main identity dimension in the sentence with _ (gap) and give the model options to select between the possible identities. We provide details of the procedure and examples in appendix C.2. However, there are many instances where a sentence is not equally probable for both contrasting identities. In order to curate sentences that serve our cause, which is mainly to provide equal opportunity, we manually select these sentences (refer to appendix D for details). After the curation process, we are left with 2416 pairs for gender and 1535 pairs for religion.
@@ 4.1 Model Selection
For our experiment we provide results for four state-of-the-art LLMs: Llama3-8b (version: Meta-Llama-3-8B-Instruct 333meta-llama/Meta-Llama-3-8B-Instruct) (AI@Meta, 2024), GPT-3.5-Turbo 444gpt-3-5-turbo, GPT-4o 555gpt-4o and Claude-3.5-Sonnet666anthropic/claude-3.5-sonnet. Since Bangla is a low resource language, not many models could generate the expected response we required. We also experimented with some other models (mentioned in limitations) but none could produce any presentable result serving our purpose.
@@ 4.2 Prompt
We design prompts to best fit the generation task. In the case of template based probing, we prompt the model for gendered role or religious identity selection, and in the case of naturally sourced probing, we use fill in the blanks approach. We provide examples of prompt creation in appendix G. In table 1, we provide the number of unique prompts for each model.
@@
Template Based Gender 2128
Religion 644
Naturally Sourced Gender 2416
Religion 1535
@@
During evaluation, the options provided to LLMs inside a prompt are randomly shuffled for both gender and religious entities to avoid selection bias Zheng et al. (2024).
@@ 4.3 Evaluation Setup
Evaluation Metric: We employ the widely used fairness metric, Disparate Impact (DI) Feldman et al. (2015), calculated as $\frac{P(Y=1|S\neq 1)}{P(Y=1|S=1)}$. For our binary identifiers (e.g., male-female, Hindu-Muslim), DI can be applied through empirical estimation. In task Q, for category a with outcomes x and y, DI is calculated as:
@@
We use occurrence frequency instead of probability and adjust the metric to adjust equal proportionality in bias scores (further justification and details is provided in appendix B):
$Bias\>Score=DI_{Q}(a)=\tanh\left(\log\frac{C_{x}(a)}{C_{y}(a)}\right)$
Here, $C_{z}$ represents the frequency of class z. We compute $DI_{G}$ and $DI_{R}$ for gender and religion biases, where $(x=female,y=male)$ and $(x=Hindu,y=Muslim)$. For a fair LLM, the DI score should be close to 0.
@@ 5 Results and Evaluation
(a) Bias Scores for Gender Bias (Positive Traits)
(b) Bias Scores for Gender Bias (Negative Traits)
(c) Bias Scores for Religious Bias (Positive Traits)
(d) Bias Scores for Religious Bias (Negative Traits)
Figure 2: Bias Scores in role selection for multiple LLMs in the case of template based probing for gender and religion data. Positive and negative traits results are shown separately. The neutral line $(Bias\>Score=0)$ is highlighted in all the figures. (Note that the results for Occupation are the same for positive and negative traits and only included in contrasting graphs for the ease of comprehending the effect of inter-mixing with other traits.)
@@ 5.1 Template Based Probing Results
We present the template based results in figure 2. We report the results based on seven different categories and include the results for positive and negative traits separately for more nuanced variations.
Gender Bias: Our findings (Figure 2(a), 2(b)) show that GPT-3.5-Turbo is consistently biased toward females, while Llama-3 and Claude-3.5-Sonnet are biased toward males across both positive and negative traits. GPT-4o exhibits the most fluctuation, switching its bias depending on the category. When the traits change from positive to negative, GPT-4o changes substantially from female direction to male direction for Personality and Communal based traits. Except for GPT-3.5-Turbo, all models display a strong male bias for occupations.
Inclusion of occupation in prompts had the most significant impact on GPT-4o, reversing its bias direction. In most other cases, occupations shifted bias scores further towards males, suggesting that LLMs place significant weight on occupation when inferring gender. High negative bias scores of Claude-3.5-Sonnet, compared to other models, may be due to the limitations in understanding Bangla context, warranting further investigation.
Religious Bias: For positive traits (Figure 2(c)), all the LLMs exhibit positive bias scores, i.e. being biased for Hindu Religion followers. All LLMs show positive scores for Occupation. The responses form GPT-4o and Llama-3 hold neutral positions for Outlook, but when associated with Occupation, their position of neutrality is compromised. For Llama-3, no specific pattern is evident and high fluctuations are noticeable.
For negative traits (Figure 2(d)), GPT models tend to adopt a neutral stance when Outlook adjectives are included in prompts. We hypothesize that the models avoid offensive responses by maintaining neutrality in negative contexts. However, GPT-4o shows a significant bias towards Muslims when negative ideological elements are present, which is concerning.
5.2 Naturally Sourced Probing Results
Gender Bias: Figure 3 shows that GPT-4o has the highest bias score, indicating a significant gender disparity in its performance. GPT-3.5, with a score just above neutral, demonstrates relatively balanced results with minor disparities. Llama-3, with a negative bias score, favors the opposite gender compared to GPT-4o but is closer to the fairness threshold. Claude-3.5-Sonnet exhibits moderate bias toward males. Notably, these scores are considerably lower than those from template-based probing.
@@ 5.2 Naturally Sourced Probing Results
Religious Bias: The bias scores for religion in Figure 3 are comparatively closer among all models. GPT-4o and Llama-3 both exhibit negative bias scores, suggesting some level of bias towards Muslims. GPT-4o exhibits the highest level of bias.
We hypothesize that, the reason for not showing substantial bias in naturally probed examples can be attributed to two points: (1) When Bangla prompt with bigger context is used from a naturally set scenario, the focus of the LLMs are not pin pointed towards the personification of a character into a religious or gender entity. (2) The guard-rails used in LLMs work better in a natural probing setting.
@@ 6 Conclusion
To summarize, by conducting experiments using two different probing techniques and dataset, we investigate gender and religion bias in multilingual LLMs in the context of Bangla. Our work demonstrates the existence of bias in both categories on different degrees. This emphasizes the need for de-biasing techniques to be applied for the use of LLMs in sensitive tasks in the realm of Bangla Language and developing proper linguistics and culturally aware framework for bias measurement. In future, we plan to investigate the effects of bias in downstream applications of Bangla language models.
@@
Our study utilized closed-source models like GPT-3.5-Turbo, GPT-4o and Claude-3.5-Sonnet which present reproducibility challenges as they can be updated at any time, potentially altering responses regardless of temperature or top-p settings. We also attempted to conduct experiments with other state-of-the-art models such as Mistral-7b-Instruct 777mistralai/Mistral-7B-Instruct-v0.2 (Jiang et al., 2023), Llama-2-7b-chat-hf 888meta-llama/Llama-2-7b-chat-hf (Touvron et al., 2023), and OdiaGenAI-BanglaLlama  [^9] OdiaGenAI/odiagenAI-bengali-base-model-v1 (Parida et al., 2023). However, these efforts were hindered by frequent hallucinations and an inability to produce coherent and presentable results. This issue underscores a broader challenge: the current limitations of LLMs in processing Bangla, a low-resource language, indicating a need for more focused development and training on Bangla-specific datasets.
@@
We also acknowledge that our results may vary with different prompt templates and datasets, constraining the generalizability of our findings. Stereotypes are likely to differ based on the context of the input and instructions. Finally our techniques all utilizes binary identities(male-female, Hindu-Muslim) for the constraints on dataset and techniques used (Please refer to appendix A). Despite these limitations, we believe our study provides essential groundwork for further exploration of social stereotypes in the context of Bangla for LLMs.
@@
Appendix A Frequency Analysis of Gender and Religion Terms in Two Bangla Corpora
We have kept our studies limited to binary genders and the major religions in Bangla speaking regions. In this section, we provide a quantitative analysis of two major Bangla corpora regarding the frequency distribution of gender and religion realted entities. We show the results in Figure 4.
We extracted the gender and religion related entities from two large corpora, BnWiki [^10] The latest bangla wiki dump used from https://dumps.wikimedia.org/bnwiki/20240901/ and Bangla2B+ Bhattacharjee et al. (2022). It is evident that there is a significant absence of non-binary genders in Bangla. For the male and female words, we used the most common male and female terms in Bangla and later aggregated the results under Men and Women terms in the data showed. The word percentages for transgenders and homosexuals are less than 2%. Note that, we used the term Hijra [^11] https://en.wikipedia.org/wiki/Hijra_(South_Asia) as an umbrella term for non-binary genders, as this semantics is prevalent in South Asia.
Figure 4: Frequency Analysis of Gender and Religious Identities in two large Bangla corpora: BnWiki and Bangla2B+
For the religion related terms, we composed the common religious identity based words in Bangla speaking regions and accommodated for their variations. In both the corpora, we can see that Hindu and Muslim related religious identities comprise of more than 70% of the total identities. Hence considering the availability of dataset, our probing techniques and corpus frequency distribution, we limited our study to binary genders and most common religions.
Appendix B Evaluation Metric Justification
Various metrics have been proposed to evaluate the fairness of LLMs. Disparate Impact compares the proportion of favorable outcomes for a minority group to a majority group, while Statistical Parity compares the percentage of favorable outcomes for monitored groups to reference groups. Metrics such as Equalized Opportunity and Equalized Odds considers ground truth. Since our dataset contains no ground truth, we chose Disparate Impact to evaluate the model responses for binary identities.
In task Q, for category a with outcomes x and y, DI is calculated as:
$DI_{Q}(a)=\frac{P(Q=x|a)}{P(Q=y|a)}$
Since we do not have probability distributions in our case, we use the occurrence frequency of each category instead. However, plotting the graphs with the above formula can be challenging because the values lie in the interval $[0,+\infty)$ with the center line in 1. For an LLM, $DI_{Q}(a)=1$ signifies perfect fairness, while values approaching 00 or +$\infty$+ ∞ indicate extreme bias towards one identity. For example, if $P(Q=female|Gender)=0.01$ and $P(Q=male|Gender)=0.99$, then $DI_{Gender}=\frac{0.01}{0.99}=0.01010101$. Conversely, if $P(Q=female|Gender)=0.99$ and $P(Q=male|Gender)=0.01$, then $DI_{Gender}=\frac{0.99}{0.01}=99$. Though both results reflect significant bias, visually interpreting these results on a graph can be difficult due to the disproportionate scaling.
To address this, we modified the metric as follows:
$Bias\>Score=DI_{Q}(a)=\tanh\left(\log\frac{C_{x}(a)}{C_{y}(a)}\right)$
Here, $C_{z}$ represents the frequency of class z. By applying the logarithmic function, we scale the values proportionally for better interpretation, and we utilize the $\tanh$ function to normalize the bias scores within the interval $[-1,1]$. A $Bias\>Score$ close to 0 indicates fairness, whereas values closer to $-1$ or $1$ indicates extreme bias towards one group or the other.
Appendix C Data Preparation
C.1 Preparing Templates